The information below pertains to Pearl Company for 2021. Net income for the year $1,250,000 7% convertible bonds issued at par ($1,000 per bond); each bond is convertible into 30 shares of common stock 2,080,000 6% convertible, cumulative preferred stock, $100 par value; each share is convertible into 3 shares of common stock 3,860,000 Common stock, $10 par value 5,710,000 Tax rate for 2021 20% Average market price of common stock $25 per share There were no changes during 2021 in the number of common shares, preferred shares, or convertible bonds outstanding. There is no treasury stock. The company also has common stock options (granted in a prior year) to purchase 75,500 shares of common stock at $20 per share.

(a) Compute basic earnings per share for 2021.

(b) Compute diluted earnings per share for 2021.
Answer #1

Answer

(a) Computation of basic earnings per share for 2021

Calculation of net income attributable to equity shareholders

= net income for the year - preference dividend

= $1250000-$231600

=$1018400

No of equity shares = 571000

Basic earnings per share = $1018400/571000 = $1.78

(b) Computation of diluted earnings per share for 2021

Calculation of adjusted profit

= net income attributable to equity shareholders + interest on bonds after tax + preference dividend

=$1018400 +$116480+$231600

=$1366480

No of shares = 62400+115800+571000+75500

= 824700

Diluted earnings per share for 2021 = $1366480/824700 = $1.66